# Low-Resolution Radar-based Gesture Recognition

### v1

@@ Abstract
TODO
@@
Distributed targets, Targets discrimination, Fuzzy shell clustering, Automotive radar, Target enumeration.
@@ I Introduction
To perform radar-based HGR, several techniques have been employed, including feature extraction followed by machine learning for classification [12, 13] and principal component analysis [14, 15]. Deep learning algorithms have significantly enhanced the performance of radar-based HGR by extracting Range-Doppler frames from FMCW radar and then using 2D CNN followed by LSTM [16, 17]. Another approach extracted Range-Doppler frames from pulse radar and used a classifier based on a TCN network [18]. Additionally, some research has focused on enhancing features from Range-Doppler maps generated from mmWave SIMO radar by using long recurrent all-convolutions neural networks (LRACN) [20], and by applying Clutter Extraction and CFAR algorithms on the Range-Doppler maps [19]. Other feature extraction methods include trajectory images from UWB-radar with 2D CNN [21]. Besides different feature extraction techniques and neural network architectures, there has also been research on different learning techniques, such as meta-learning with custom loss on mmWave radar [22] or enhancing HGR accuracy by adding micro-Doppler information [23, 24, 25].
The effectiveness of radar-based gesture recognition critically hinges on two key aspects of radar resolution: range resolution and Doppler resolution. Previous works focus on high classification accuracy and robustness, selecting radar resolution according to classification task needs, thus balancing low-cost, low-resolution radar with classification performance. Super-resolution techniques can resolve this trade-off by replacing high-cost, high-resolution radar with low-cost, low-resolution radar enhanced by super-resolution techniques in space and frequency. There is extensive work on radar-based super-resolution techniques like classical direction of arrival (DOA) estimation, such as MUSIC and ESPRIT [26, 27, 28], but these methods require non-coherent sources, multiple receivers, fewer sources than receivers, and many samples for good performance. To make these methods more robust and applicable in real life, DOA-based deep learning has been developed [29, 30, 31, 32, 33], but assumptions do not hold in the case of HGR, such as the number of sources being equal to or smaller than the number of sensors. Additionally, this approach focuses on angular separation and not on distance and Doppler frequency separation.
@@
    The networks are trained simultaneously using a combination of super-resolution and classification losses, balanced as a trade-off between the two.
@@ II Problem Definition
$\begin{split}f(t,T)=p(t-nT_{P})e^{j2\pi f_{c}(t-nT_{p})}\forall n=0,\ldots,M-1\;,\end{split}$ (1)
@@ III The Proposed Approach
This section presents and explains the pipeline of our method, depicted in Figure 2. First, the low-resolution data $D_{lr}$ undergoes pre-processing in the ”LR Pre-Processing” block before entering to SuperGestNet block. SuperGestNet block that illustrate in Figure 3 consists of three sub-blocks.
The first sub-block is the ”Super-Resolution Model” block, which takes $D_{lr}$ as input and outputs $D_{sr}$, next is the ”Classification Pre-Processing” block, which processes $D_{sr}$ and outputs $\hat{D}_{sr}$ and finally, the ”Classifier Model” block uses $\hat{D}_{sr}$ to make the classification decision. The final block is the ”Training Loss” block, which describes our combined SR and classification training loss.
@@ III-A LR Pre-Processing
    Complex to Real-Imaginary Transformation: The complex-valued data is split into its real and imaginary components. This step is performed by adding a new channel to the data cube $R$, resulting in an output shape of $(C,F,\frac{M}{d_{s}},\frac{N_{f}}{d})$, where $C$ represents the two channels corresponding to the real and imaginary parts of the signal.
@@ IV-C Training Details
This section provides technical details on the training process. We trained our model, SuperGestNet, denoted by $M_{\gamma}^{d}$, using different combinations of $\gamma$ and $d$, where $d=d_{f}=d_{d}$. The parameter $\gamma$, representing the weight of the super-resolution training loss, was set to 0, 0.5, 1, and 2. The parameter $d$, the down-sampling factor, was set to 2, 4, and 8. When $d=1$, no down-sampling was performed.
We also trained our classifier with the same down-sampling factors, using cubic interpolation instead of a deep learning model, where $\gamma$ is not applicable. This baseline is denoted by $c$. Additionally, we trained the super-resolution model SAFMN on the data, then trained the classifier with the super-resolved data from the SAFMN model, denoted by $M_{\gamma}^{d}$ with $\gamma$ represented by ’*’. Both of these models serve as benchmarks in our experiments.
Note, that a different model is trained for each set of $d$ and $\gamma$ parameters.
For the super-resolution model, we set the number of channels in the first 3x3 convolution layer to 36, so each scale level in the SAFM layer has 9 channels, we used 8 FFM blocks and the up-scaling factor set according to $d$ along both fast and slow axes.
@@ IV-D Down-sample Effect on Classification Accuracy
To understand how lower resolution impacts classification accuracy in both distance (fast time axis) and Doppler (slow time axis), we trained $M_{d}$ on lower resolution data. First, we examined the effect of Doppler resolution on classifier performance by setting $d_{f}=1$ and varying $d_{s}$. The decrease in accuracy is shown in Figure 8. Next, we assessed the impact of distance resolution by setting $d_{s}=1$ and varying $d_{s}$. The decrease in accuracy is shown in Figure 9. As can be seen, the Doppler resolution has a greater impact on classification accuracy compared to distance resolution, this indicates that features from the Doppler axis are more critical for classification. Additionally, the results emphasize the sensitivity of the classification to the resolution of the data and the need to develop a super resolution method that can highlight the features relevant to the classification.
@@ IV-E Performance of the Super-Resolution
The purpose of super-resolution (SR) is to extract important features to enhance classification accuracy. Each model uses a different approach for SR. TinyRadarNN employs traditional ’cubic’ interpolation, SafmnTiny uses a pre-trained SR model (SAFMN), and SuperGestNet is trained to perform SR and classification together as a cascade model.
In the first two models, the SR task is performed separately from the classification task, so classification does not impact the SR results. In contrast, SuperGestNet is trained with a training loss that balances between SR and classification tasks. The classification loss can be considered a constraint that guides SR to enhance relevant features for classification accuracy. Table II present the results of the super-resolution metrics of $M_{\gamma}^{d}$ under different settings of $\gamma$ and $d$. For any value of $d$ the model $M_{*}^{d}$ outperforms $M_{\gamma}^{d}$ for $\gamma$ = 0, 0.5, 1, 2 on the super-resolution metrics. This is expected because SafmnTiny uses an SR model trained solely on the SR task, while SuperGestNet is trained for SR with a classification constraint. The effect of $\gamma$ on the SR result is clear; a higher $\gamma$ means higher priority for the SR task, which leads to better performance in the SR task.
An example of SR is shown in Figure 10. First, we apply SR to the data, then perform FFT on the fast axis and take the absolute value to obtain the range-Doppler map. The result of SafmnTiny is closer to the original data compared to SuperGestNet.
(a) No down-sampling
(b) Down-sampling factor of 4
(c) SR model alone
(d) SR from SR classifier
Figure 10: Abs Range-Doppler maps with different down-sampling factors: (a) no down-sampling, (b) down-sampling factor of 4, (c) SR model alone, (d) SR from SR classifier.
@@ IV-F Effect of Gamma on Classification Accuracy
Table II presents the classification accuracy of TinyRadarNN, SafmnTiny, and SuperGestNet under various down-sampling factors and $\gamma$ values. TinyRadarNN and SafmnTiny serve as benchmark models to assess the impact of Super-Resolution (SR) on improving classification accuracy across down-sampling factors $d_{s}=1,2,4,8$ and to highlight the difference between jointly and separately training the super-resolution model and the classifier. The primary focus is on classification accuracy, which is crucial for evaluating the effectiveness of each model.
@@
For $d=1$, TinyRadarNN achieves an accuracy of 83.95%. This serves as the baseline for comparison when no down-sampling is applied.
@@
When $d=2$, the accuracy results show variations among different models. TinyRadarNN’s accuracy slightly decreases to 81.85%. SafmnTiny, which incorporates a pre-trained SR model, achieves an accuracy of 83.16%, indicating its capability to recover important features lost due to down-sampling. SuperGestNet models exhibit a range of accuracies depending on the $\gamma$ value. SuperGestNet2 achieves 84.00%, while the best performance is observed with SuperGestNet1, reaching an accuracy of 85.20%. This suggests that balancing SR and classification tasks (with $\gamma=1$) can improve accuracy. However, when $\gamma$ is reduced further, SuperGestNet0.5 still performs well at 84.80%, while SuperGestNet0 shows a significant drop to 82.28%, highlighting the importance of SR in enhancing classification accuracy.
@@
As $d$ increases to 4, the accuracy of all models decreases, reflecting the greater challenge posed by higher down-sampling. TinyRadarNN’s accuracy drops to 77.93%. SafmnTiny maintains relatively high accuracy at 79.16%. SuperGestNet models show a similar trend to $d=2$, with SuperGestNet0.5 achieving the highest accuracy at 81.46%, followed by SuperGestNet2 and SuperGestNet1 with 81.34% and 80.91%, respectively. SuperGestNet0 again performs poorly with an accuracy of 78.20%.
@@
For $d=8$, the accuracy further declines due to the increased difficulty of the task. TinyRadarNN reaches 74.10%. SuperGestNet models show varying performance with the highest accuracy from SuperGestNet0.5 at 76.78%. SuperGestNet2 and SuperGestNet1 achieve 76.62% and 76.04%, respectively. The performance of SuperGestNet0 drops significantly to 73.35%, indicating the necessity of SR for maintaining classification accuracy at higher down-sampling levels.
@@
In summary, for each down-sampling factor, SuperGestNet outperforms SafmnTiny in classification accuracy, although SafmnTiny achieves better performance on the SR task. This emphasizes the advantage of joint training of the tasks, with classification used as regularization for SR. Additionally, we see a dependency of the $\gamma$ value on the down-sampling factor. As the down-sampling factor increases, $\gamma$ decreases. This is because the corruption of the data is greater when the down-sampling factor increases, and SR cannot completely recover all the missing data. In such cases, a smaller $\gamma$ will give less weight to the SR task.
IV-G Training Networks Separately
To demonstrate the contribution of our custom training loss function, which enforces the super-resolution model to enhance relevant features for classification, we first trained the super-resolution model without the classification model using the $L_{1}$ loss function. In a subsequent training, we used the pre-trained super-resolution model, froze its weights, and trained only the classifier using the CrossEntropy loss function. The input for the classifier was the super-resolution data from our frozen super-resolution network. compeer to the cascade model that achieve accuracy of 81.46 this model only achieve 79.24
IV-H 3 SR Models and 1 Classifier
The purpose of the Super-Resolution (SR) model in SuperGestNet is to enhance the relevant features for classification. The input dimensions for the classifier remain constant, and the performance metrics of the super-resolution are similar for down-sampling factors of 2 and 4, as shown in Table III. This similarity indicates that the difference between the super-resolved data for down-sampling factors of 2 and 4 is not significant.
Given these observations, we propose MultiSuperGestNet, which incorporates three SR models corresponding to down-sampling factors of 2, 3, and 4, along with a single classifier. To train MultiSuperGestNet, we utilized our dataset and, for each batch, sampled the data with different down-sampling factors. For each batch, if the data was sampled with a down-sampling factor of 2, the SR model for factor 2 was chosen; if the data was sampled with a down-sampling factor of 3, the SR model for factor 3 was chosen; similarly, for a down-sampling factor of 4, the SR model for factor 4 was chosen. This training method effectively increased our dataset size by a factor of three because each data sample produced three training samples—one for each down-sampling factor.
The results from Table III show that the classification accuracy increased for each down-sampling factor. This makes sense because we have a single classifier trained on a dataset that is effectively three times larger. There is a small decrease in the SR metrics, but it is not significant.
@@ IV-G Training Networks Separately
Table II: Performance Metrics for OurModel with different down-sampling factors d and different $\gamma$ values
@@
TinyRadarNN 81.85 0.215 0.756 14.93
SafmnTiny 83.16 0.033 0.936 21.7
SuperGestNet2 84.00 0.065 0.871 19.18
SuperGestNet1 85.20 0.074 0.852 18.4
SuperGestNet0.5 84.80 0.093 0.817 16.98
SuperGestNet0 82.28 1.167 0.342 -3.45
@@
TinyRadarNN 77.93 0.274 0.711 13.87
SafmnTiny 79.16 0.061 0.876 19.93
SuperGestNet2 81.34 0.085 0.857 17.46
SuperGestNet1 80.91 0.093 0.847 16.93
SuperGestNet0.5 81.46 0.108 0.833 15.77
SuperGestNet0 78.20 1.355 0.294 -5.54
@@
TinyRadarNN 74.10 - - -
SuperGestNet2 76.62 0.117 0.826 15.48
SuperGestNet1 76.04 0.134 0.818 14.36
SuperGestNet0.5 76.78 0.167 0.794 12.43
SuperGestNet0 73.35 0.45 0.578 -10.96

### v2

@@ Abstract
This article addresses the challenge of recognizing hand gestures using low-resolution radar. Conventional sensors like cameras, infrared, and ultrasonic have limitations in lighting and range, while radars can provide more reliable performance. To leverage radar in gesture recognition, an innovative approach, SuperGestNet, which combines super-resolution and gesture classification networks trained jointly, is proposed. SuperGestNet consists of three sub-components: a super-resolution model that enhances radar data resolution, a pre-processing step for gesture classification, and a classification model optimized for low-power devices. Innovative pre-processing and data augmentation techniques are introduced to transform low-resolution radar data into high-resolution images using a spatially adaptive feature modulation network. This approach aims to enhance classification accuracy by generating high-quality radar images that emphasize critical features for gesture recognition. The model demonstrates improvements in both super-resolution and classification accuracy over traditional methods, making it a promising solution for reliable, low-power radar-based gesture recognition in real-world applications.
@@
Low-resolution radar, Gesture classification, Hand gesture classification, Super-resolution
@@ I Introduction
To perform radar-based HGR, several techniques have been employed, including feature extraction followed by machine learning for classification [12, 13] and principal component analysis (PCA). PCA has been utilized to reduce the dimensionality of large datasets, such as Gray Level Co-occurrence Matrix (GLCM) texture features extracted from EEG spectrogram images for IQ classification [14]. In another application, PCA was used to generate radiation patterns for a microwave compressive sensing system, optimizing the measurement process by requiring fewer samples compared to random-pattern-based methods [15]. Deep learning algorithms have significantly enhanced the performance of radar-based HGR by extracting Range-Doppler frames from FMCW radar and then using 2D CNN followed by LSTM [16, 17]. Alternatively in [18] Range-Doppler frames from pulse radar are extarcted and a classifier based on a TCN network is used. Additionally, some research has focused on enhancing features from Range-Doppler maps generated from mmWave SIMO radar by using long recurrent all-convolutions neural networks (LRACN) [20], and by applying Clutter Extraction and CFAR algorithms on the Range-Doppler maps [19]. Other feature extraction methods include trajectory images from UWB-radar with 2D CNN [21]. Besides different feature extraction techniques and neural network architectures, there has also been research on different learning techniques, such as meta-learning with custom loss on mmWave radar [22] or enhancing HGR accuracy by adding micro-Doppler information [23, 24, 25].
The effectiveness of radar-based gesture recognition critically hinges on two types of radar resolution: range resolution and Doppler resolution. Previous works focus on high classification accuracy and robustness, selecting radar resolution according to classification task needs, thus adjusting low-cost, low-resolution radar to required classification performance. Super-resolution techniques can resolve this trade-off by replacing high-cost, high-resolution radar with low-cost, low-resolution radar enhanced by super-resolution techniques in space and frequency. There is extensive work on radar-based super-resolution techniques like classical direction of arrival (DOA) estimation, such as MUSIC and ESPRIT [26, 27, 28], but these methods require non-coherent sources, multiple receivers, fewer sources than receivers, and many samples for good performance. In order to make these methods more robust and applicable in real life, DOA-based deep learning has been developed [29, 30, 31, 32, 33]. However these assumptions do not hold in the case of HGR, such as the number of sources being equal to or smaller than the number of sensors. Additionally, this approach focuses on angular separation and not on distance and Doppler frequency separation.
@@
    The networks are trained jointly using a weighed combination of super-resolution and classification losses, thus managing a trade-off between the two tasks.
@@
  4. 4.
    New training is explored enabling usage of shared classification network independent of the data resolution
  5. 5.
    We propose and analyze a novel recursive architecture for super-resolution motivated by multi-scale theory.
@@ II Problem Definition
$\begin{split}f(t,T)=p(t-nT_{P})e^{j2\pi f_{c}(t-nT_{p})},\forall n=0,\ldots,M-1\;,\end{split}$ (1)
@@ III The Proposed Approach
This section presents and explains the pipeline of our method, depicted in Figure 2. First, the low-resolution data $D_{lr}$ undergoes pre-processing in the ”LR Pre-Processing” block before entering to SuperGestNet block. SuperGestNet block illustrated in Figure 3 consists of three sub-blocks.
The first sub-block is the ”Super-Resolution Model” block, which takes $D_{lr}$ as input and outputs $D_{sr}$, followed by the ”Classification Pre-Processing” block, which processes $D_{sr}$ and outputs $\hat{D}_{sr}$ and finally, the ”Classifier Model” block uses $\hat{D}_{sr}$ to perform classification.
@@ III-A LR Pre-Processing
    Complex to Real-Imaginary Transformation: The complex-valued data is split into its real and imaginary components. This step is performed by adding a new channel to the data cube $R$, resulting in an output shape of $(C,F,\frac{M}{d_{s}},\frac{N}{d_{f}})$, where $C$ represents the two channels corresponding to the real and imaginary parts of the signal.
@@ IV-C Training Details
This section provides technical details on the training process. We trained our model, SuperGestNet, denoted by $M_{\gamma}^{d}$, using different combinations of $\gamma$ and $d$, where $d=d_{f}=d_{s}$. The parameter $\gamma$, representing the weight of the super-resolution training loss, was set to 0, 0.5, 1, and 2. The parameter $d$, the down-sampling factor, was set to 2, 4, and 8. When $d=1$, no down-sampling was performed.
We also trained our classifier with the same down-sampling factors, using cubic interpolation instead of a deep learning model, where this model name is CubicTinyRadarNN and denoted by $C^{d}$. Additionally, we trained the super-resolution model SAFMN on the data, then froze SAFMN and trained the classifier with the super-resolved data from the SAFMN model,this model name is SafmnTiny and denoted by $FM^{d}$. Both of these models serve as benchmarks in our experiments.
Note that a different model is trained for each set of $d$ and $\gamma$ parameters.
For the super-resolution (SAFMN) model, we set the number of channels in the first $3\times 3$ convolution layer to 36, so each scale level in the SAFM layer has 9 channels. We used 8 FFM blocks, and the up-scaling factor was set according to $d$ along both fast and slow axes.
Next, we tried to improve the classification accuracy by using 3 different SAFMN models, each one trained for a different down-sampling factor $d$, and one classifier for all three super-resolution models. This model is named MultiSuperGestNet and denoted by $SM$.
We then used SAFMN as a recursive super-resolution model, applying the model iteratively to achieve super-resolution by a factor of 2 each time. To up-sample the image by a factor of 4, we applied SAFMN twice before classification. This model is named RecSuperGestNet and denoted by $RM$.
For RecSuperGestNet and MultiSuperGestNet we used $\gamma=2$ and the same parameters for SAFMN like before.
@@ IV-D Down-sample Effect on Classification Accuracy
To understand how lower resolution impacts classification accuracy in both distance (fast time axis) and Doppler (slow time axis), we trained $C_{d}$ on lower resolution data. First, we examined the effect of Doppler resolution on classifier performance by setting $d_{f}=1$ and varying $d_{s}$. The decrease in accuracy is shown in Figure 8. Next, we assessed the impact of distance resolution by setting $d_{s}=1$ and varying $d_{s}$. The decrease in accuracy is shown in Figure 9. As can be seen, the Doppler resolution has a greater impact on classification accuracy compared to distance resolution, this indicates that features from the Doppler axis are more critical for classification. Additionally, the results emphasize the sensitivity of the classification to the resolution of the data and the need to develop a super resolution method that can highlight the features relevant to the classification.
@@ IV-E Performance of the Super-Resolution
The purpose of super-resolution (SR) is to extract important features to enhance classification accuracy. Each model uses a different approach for SR. CubitTinyRadarNN ($C^{d}$) employs traditional ’cubic’ interpolation, SafmnTiny ($FM_{d}$) uses a pre-trained SR model (SAFMN), and SuperGestNet ($M_{d}^{\gamma}$) is trained to perform SR and classification together as a cascade model.
In the first two models, the SR task is performed separately from the classification task, so classification does not impact the SR results. In contrast, $M_{d}^{\gamma}$ is trained with a training loss that balances between SR and classification tasks. The classification loss can be considered a constraint that guides SR model to enhance relevant features for classification accuracy. Table II present the results of the super-resolution metrics of $M_{\gamma}^{d}$, $FM^{d}$ , $C^{d}$ under different settings of $\gamma$ and $d$. For any value of $d$ the model $FM^{d}$ outperforms $M_{\gamma}^{d}$ for all the $\gamma$ values on the super-resolution metrics. This is expected because $FM^{d}$ uses an SR model trained solely on the SR task, while $M_{\gamma}^{d}$ is trained for SR with a classification constraint. The effect of $\gamma$ on the SR result is clear, a higher $\gamma$ means higher priority for the SR task, which leads to better performance in the SR task.
An example of SR is shown in Figure 10. First, we apply SR to the data, then perform FFT on the fast axis and take the absolute value to obtain the range-Doppler map. The result of $FM^{4}$ is closer to the original data compared to $M^{4}_{2}$.
Figure 10: Abs Range-Doppler maps (a) original high resolution data, (b) doopler map that generated from down sample data by factor of 4, (c) super-resolution doopler map that generated with $FM^{4}$, (d) super-resolution doopler map that generated with $M^{4}_{2}$.
@@ IV-F Effect of Gamma on Classification Accuracy
Table II presents the classification accuracy of $C^{d}$, $FM^{d}$, and $M_{\gamma}^{d}$ under various down-sampling factors $d$ and $\gamma$ values. $C^{d}$ and $FM^{d}$ serve as benchmark models to assess the impact of Super-Resolution (SR) on improving classification accuracy across down-sampling factors $d=d_{f}=d_{s}=1,2,4,8$ and to highlight the difference between jointly and separately training the super-resolution model and the classifier. The primary focus is on classification accuracy, which is crucial for evaluating the effectiveness of each model.
@@
For $d=1$, $C^{1}$ achieves an accuracy of 83.95%. This serves as the baseline for comparison when no down-sampling is applied.
@@
When $d=2$, the accuracy results show variations among different models. $C^{2}$ accuracy slightly decreases to 81.85%. $FM^{2}$, which incorporates a pre-trained SR model, achieves an accuracy of 83.16%, indicating its capability to recover important features lost due to down-sampling. $M_{\gamma}^{2}$ models exhibit a range of accuracies depending on the $\gamma$ value. $M_{2}^{2}$ achieves 84.00%, while the best performance is observed with $M_{1}^{2}$, reaching an accuracy of 85.20%. This suggests that balancing SR and classification tasks (with $\gamma=1$) can improve accuracy. However, when $\gamma$ is reduced further, $M_{0.5}^{2}$ still performs well at 84.80%, while $M_{0}^{2}$ shows a significant drop to 82.28%, highlighting the importance of SR in enhancing classification accuracy.
@@
As $d=4$ increases to 4, the accuracy of all models decreases, reflecting the greater challenge posed by higher down-sampling. $C^{4}$ accuracy drops to 77.93%. $FM^{4}$ maintains relatively high accuracy at 79.16%. $M_{\gamma}^{4}$ models show a similar trend to $d=2$, with $M_{0.5}^{4}$ achieving the highest accuracy at 81.46%, followed by $M_{2}^{4}$ and $M_{1}^{4}$ with 81.34% and 80.91%, respectively. $M_{0}^{4}$ again performs poorly with an accuracy of 78.20%.
@@
For $d=8$ , the accuracy further declines due to the increased difficulty of the task. $C^{8}$ reaches 74.10%. $M_{\gamma}^{8}$ models show varying performance with the highest accuracy from $M_{0.5}^{8}$ at 76.78%. $M_{2}^{8}$ and $M_{1}^{8}$ achieve 76.62% and 76.04%, respectively. The performance of $M_{0}^{8}$ drops significantly to 73.35%, indicating the necessity of SR for maintaining classification accuracy at higher down-sampling levels.
@@
In summary, for each down-sampling factor, SuperGestNet ($M_{\gamma}^{d}$) outperforms SafmnTiny ( $FM^{d}$) in classification accuracy, although $FM^{d}$ achieves better performance on the SR task. This emphasizes the advantage of joint training of the tasks, with classification used as regularization for SR. Additionally, we see a dependency of the $\gamma$ value on the down-sampling factor. As the down-sampling factor increases, $\gamma$ decreases. This is because the corruption of the data is greater when the down-sampling factor increases, and SR cannot completely recover all the missing data. In such cases, a smaller $\gamma$ will give less weight to the SR task.
@@
Table II: Performance Metrics for CubitTinyRadarNN, SafmnTiny and SuperGestNet with different down-sampling factors d and different $\gamma$ values
@@
$C^{2}$ 81.85 0.215 0.756 14.93
$FM^{2}$ 83.16 0.033 0.936 21.7
$M_{2}^{2}$ 84.00 0.065 0.871 19.18
$M_{1}^{2}$ 85.20 0.074 0.852 18.4
$M_{0.5}^{2}$ 84.80 0.093 0.817 16.98
$M_{0}^{2}$ 82.28 1.167 0.342 -3.45
@@
$C^{4}$ 77.93 0.274 0.711 13.87
$FM^{4}$ 79.16 0.061 0.876 19.93
$M_{2}^{4}$ 81.34 0.085 0.857 17.46
$M_{1}^{4}$ 80.91 0.093 0.847 16.93
$M_{1}^{4}$ 81.46 0.108 0.833 15.77
$M_{0}^{4}$ 78.20 1.355 0.294 -5.54
@@
$C^{8}$ 74.10 - - -
$M_{2}^{8}$ 76.62 0.117 0.826 15.48
$M_{1}^{8}$ 76.04 0.134 0.818 14.36
$M_{0.5}^{8}$ 76.78 0.167 0.794 12.43
$M_{0}^{8}$ 73.35 0.45 0.578 -10.96
IV-G Training Networks Separately
The purpose of this section is to highlight the difference between SafmnTiny and SuperGestNet. Although both models share the same architecture, the main difference lies in their training processes and the applied loss functions. To train SafmnTiny, we first trained the super-resolution (SR) model, SAFMN, using the L1 loss. After training, we froze the model and proceeded to train the classifier separately with Cross Entropy loss, utilizing the super-resolved data generated by SAFMN. These two training steps were performed independently, so neither training phase influenced the other. Training a super-resolution model typically aims to create high-resolution images that appear visually appealing and realistic. However, standard super-resolution training does not guarantee that the generated high-resolution data accurately reflects the original content. This can lead to artifacts or added features that may not have existed in the original image but are imperceptible to the human eye. In contrast, SuperGestNet employs a cascade training approach where both models are trained jointly. The combined loss function, as defined in equation (15), integrates the L1 loss for super-resolution and the Cross Entropy loss for classification. By balancing these components with the parameter $\gamma$, this approach ensures that the super-resolution model enhances only the features necessary for accurate classification while constraining it from adding non-existent features to the data. This joint training strategy aligns the super-resolution task more closely with the classification objective, leading to more relevant and reliable outputs.
In Figure 10we can observe the original Doppler map (a), the Doppler map generated from down-sampled data by a factor of 4 on both axes (b), the Doppler map produced by SafmnTiny (c), and the Doppler map generated by SuperGestNet (d). The results from SafmnTiny appear more realistic and closer to the original data compared to those from SuperGestNet, which is attributed to the distinct training processes and the absence of classification loss in SafmnTiny. Notably, most of the data in the Doppler map is zero, indicating that enhancing the entire image is unnecessary. Focusing on the relevant parts helps improve classification performance and avoids introducing modifications that do not contribute to the task.
Table II shows that the super-resolution metrics for SafmnTiny are superior to those for SuperGestNet, indicating that SafmnTiny performs better on the entire dataset. However, the classification accuracy of SuperGestNet surpasses SafmnTiny’s across all values of $d$ and $\gamma$. This suggests that while SafmnTiny excels in generating visually accurate high-resolution images, SuperGestNet’s joint training approach enhances the features relevant for classification, resulting in better performance in classification tasks.
The comparison between SafmnTiny and SuperGestNet underscores the trade-offs between separate and joint training strategies. While SafmnTiny produces superior overall image quality, SuperGestNet’s cascade approach effectively balances image quality and classification accuracy. This demonstrates that, for applications where classification performance is critical, integrating classification loss into the training process can yield better results by tailoring super-resolution enhancements to task-specific requirements.
IV-H MultiSuperGestNet Model
The purpose of the super-resolution model in SuperGestNet is to enhance the relevant features for classification. The input dimensions for the classifier remain constant, and the performance metrics of the super-resolution are similar for down-sampling factors of 2 and 4, as shown in Table III. This similarity indicates that the difference between the super-resolved data for down-sampling factors of 2 and 4 is not significant.
Given these observations, we propose MultiSuperGestNet, which incorporates three SR models corresponding to down-sampling factors of 2, 3, and 4, along with a single classifier. To train MultiSuperGestNet, we utilized our dataset and, for each batch, sampled the data with different down-sampling factors. For each batch, if the data was sampled with a down-sampling factor of 2, the SR model for factor 2 was chosen; if the data was sampled with a down-sampling factor of 3, the SR model for factor 3 was chosen; similarly, for a down-sampling factor of 4, the SR model for factor 4 was chosen and we set $\gamma=2$ for all super-resolution models. This training method effectively increased our dataset size by a factor of three because each data sample produced three training samples—one for each down-sampling factor.
The results in Table III indicate an improvement in classification accuracy for each down-sampling factor. This is expected, as the single classifier is trained on an augmented dataset that is effectively three times larger. While there is a slight decrease in the SR metrics, it is not substantial. Furthermore, MultiSuperGestNet reduces memory usage by consolidating the weights of 3 classifiers into a single model, allowing it to handle three different resolutions efficiently.
Table III presents the results for MultiSuperGestNet. While there is a slight decrease in the super-resolution metrics, an improvement in classification accuracy is observed.